OpenStudy (anonymous):

Simplify expression (3x^5/y^2)^4

OpenStudy (jdoe0001):

\(\large \bf \left(\cfrac{3x^5}{y^2}\right)^{\color{brown}{ 4}}\implies \cfrac{(3x^5)^{\color{brown}{ 4}}}{(y^2)^{\color{brown}{ 4}}}\implies \cfrac{3^{\color{brown}{ 4}}x^{5\cdot {\color{brown}{ 4}}}}{y^{2\cdot {\color{brown}{ 4}}}}\)

OpenStudy (jdoe0001):

you'd distribute the exponent

OpenStudy (anonymous):

 Do you under stand that or need more explanation?